This Question: 1 pt Find the indicated value of the give g(x) = 6x? - 9; 9(6) g(6)=
Answer #1

Q “g(x) = 6x2-9 . To find g(6) Put a=6 in the function 916): 616)2-9 916)=6.36-9 9 (6)=216-9 1g (6)= 207]